### v1

# Mapping the sources of CP violation in neutrino oscillations from the seesaw mechanism

Abstract
We present the first complete calculation of the Jarlskog invariant, a working measure of the strength of CP violation in the flavor oscillations of three light neutrino species, with the help of a full Euler-like block parametrization of the flavor structure in the canonical seesaw mechanism. We find that this invariant depends on 240 linear combinations of the 6 original phase parameters that are responsible for CP violation in the decays of three heavy Majorana neutrinos in 27 linear combinations as a whole, and thus provides the first model-independent connection between the microscopic and macroscopic matter-antimatter asymmetries.
1 Motivation
The most natural and most economical extension of the standard model (SM) of particle physics to generate tiny masses for the three active neutrinos, whose flavor and mass eigenstates are denoted respectively as $\nu_{\alpha}$ (for $\alpha=e,\mu,\tau$) and $\nu_{i}$ (for $i=1,2,3$), is to (a) introduce three right-handed neutrino fields $N_{\alpha\rm R}$ in correspondence to $\nu_{\alpha\rm L}$, (b) allow the Yukawa interactions between the SM Higgs doublet and the chiral neutrino fields, and (c) permit the harmless Majorana mass term [1] formed by the fields $N_{\alpha\rm R}$ and their charge-conjugated states $(N_{\alpha\rm R})^{c}$ [2, 3, 4, 5, 6]. This widely accepted seesaw mechanism constitutes a canonical theory for the origin of neutrino masses, and it is fully consistent with the spirit of the SM effective field theory after the relevant heavy degrees of freedom are integrated out [7]. Furthermore, the lepton-number-violating and CP-violating decays of three heavy Majorana neutrino mass eigenstates $N_{j}$ (for $j=4,5,6$) give rise to the well-known leptogenesis mechanism [8] which can help explain the observed baryon-antibaryon asymmetry of the Universe [9] in a rather natural way. So the seesaw mechanism offers a particularly appealing and convincing killing-two-birds-with-one-stone landscape for the fundamental parts of both particle physics and cosmology [10].
The aforementioned seesaw mechanism consists of 18 free parameters: 3 heavy Majorana neutrino masses $M_{j}$ (for $j=4,5,6$), 9 active-sterile flavor mixing angles $\theta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) and 6 independent CP-violating phases $\alpha_{i}$ and $\beta_{i}$ (for $i=1,2,3$), where $\theta_{ij}$, $\alpha_{i}$ and $\beta_{i}$ are the Euler-like parameters used to describe the seesaw flavor structure [11, 12]. Except $M_{i}$, which are expected to be far above the scale of electroweak symmetry breaking, the other 15 parameters all appear in the leptonic weak charged-current interactions:
$-{\cal L}_{\rm cc}=$ $\hskip 2.84544pt\frac{g}{\sqrt{2}}\hskip 2.84544pt\overline{\big{(}\begin{matrix}e&\mu&\tau\end{matrix}\big{)}_{\rm L}}\hskip 2.84544pt\gamma^{\mu}\left[U\left(\begin{matrix}\nu_{1}\cr\nu_{2}\cr\nu_{3}\cr\end{matrix}\right)_{\rm L}+R\left(\begin{matrix}N_{4}\cr N_{5}\cr N_{6}\cr\end{matrix}\right)_{\rm L}\hskip 1.42271pt\right]W^{-}_{\mu}+{\rm h.c.}\;,$ (1)
in which $U=AU_{0}$ denotes the $3\times 3$ Pontecorvo-Maki-Nakagawa-Sakata (PMNS) lepton flavor mixing matrix [13, 14, 15] with $U_{0}$ being a unitary matrix, and $R$ measures the strengths of Yukawa interactions of massive neutrinos and satisfies the unitarity condition $UU^{\dagger}+RR^{\dagger}=AA^{\dagger}+RR^{\dagger}=I$. Note that $A$ and $R$ contain the same active-sterile flavor mixing and CP violation parameters, and $U$ is also correlated with $R$ via the exact seesaw relation $UD_{\nu}U^{T}=-RD_{N}R^{T}$ [11, 12], where $D_{N}\equiv{\rm Diag}\{M_{4},M_{5},M_{6}\}$ and $D_{\nu}\equiv{\rm Diag}\{m_{1},m_{2},m_{3}\}$ with $m_{i}$ being the resulting light Majorana neutrino masses (for $i=1,2,3$). In this case the 9 light degrees of freedom of $D_{\nu}$ and $U_{0}$ (i.e., the 3 neutrino masses, 3 flavor mixing angles and 3 CP-violating phases) are derivational in the sense they can be fully determined from the 18 original seesaw flavor parameters [16],
$U_{0}D_{\nu}U^{T}_{0}=-\left(A^{-1}R\right)D_{N}\left(A^{-1}R\right)^{T}\;.$ (2)
In other words, Eq. (2) offers a feasible way to test the seesaw mechanism at low energies no matter how insignificant the non-unitarity of $U$ (i.e., the deviation of $U$ from $U_{0}$) is  [^1] One may take account of slight quantum corrections between the seesaw and electroweak scales by means of the renormalization-group equations of the relevant flavor parameters [17, 18, 19]..
The phase parameters hidden in $A^{-1}R$ are responsible for CP violation in three heavy Majorana neutrino decays at high energies and must manifest themselves in the flavor oscillations of three light neutrinos such as $\nu_{\mu}\to\nu_{e}$ and $\overline{\nu}_{\mu}\to\overline{\nu}_{e}$. It is the unique Jarlskog invariant [20, 21]  [^2] To avoid any possible confusion, we use $(i,i^{\prime},i^{\prime\prime})$ to denote the light neutrino indices $(1,2,3)$, and $(j,k,l)$ to denote the heavy neutrino indices $(4,5,6)$ throughout this paper.
${\cal J}_{\nu}\equiv{\rm Im}\left[\left(U_{0}\right)_{\alpha i}\left(U_{0}\right)_{\beta i^{\prime}}\left(U_{0}\right)^{*}_{\alpha i^{\prime}}\left(U_{0}\right)^{*}_{\beta i}\right]\;,$ (3)
in which the Greek subscripts $(\alpha,\beta)$ run cyclically over $(e,\mu,\tau)$ and the Latin subscripts $(i,i^{\prime})$ run cyclically over $(1,2,3)$, that universally measures the strength of CP-violating effects in neutrino oscillations in the $U\to U_{0}$ limit. Given the fact that $U\simeq U_{0}$ is an excellent approximation (i.e., the non-unitarity of $U$ has been well constrained from a global analysis of current data on flavor and electroweak precision measurements [22, 23, 24, 25]), ${\cal J}_{\nu}$ is expected to be a working description of leptonic CP violation at low energies to a good degree of accuracy. The T2K long-baseline neutrino oscillation experiment has so far ruled out ${\cal J}_{\nu}=0$ at the $2\sigma$ level [26, 27].
So a model-independent calculation of ${\cal J}_{\nu}$ in terms of the 6 original CP-violating phases in the seesaw mechanism is fundamentally important to establish a testable correlation between the microscopic and macroscopic matter-antimatter asymmetries. But such a calculation has been lacking, as a general and direct connection between the effects of CP violation at high and low energies was concluded to be impossible in the seesaw framework due to its unknown flavor structure [28].
We are going to solve this long-standing problem by presenting the first complete calculation of the Jarlskog invariant ${\cal J}_{\nu}$ with the help of a full Euler-like block parametrization of the seesaw flavor texture proposed in Refs. [11, 12]. We demonstrate that ${\cal J}_{\nu}$ depends directly on 240 linear combinations of the 6 original phase parameters in $R$, which are responsible for CP violation in the lepton-number-violating decays of three heavy Majorana neutrinos in 27 linear combinations as a whole and thus for the validity of leptogenesis. This model-independent result will help open an important bottom-up window to probe the cosmic matter-antimatter asymmetry from CP violation in neutrino oscillations at low energies via the canonical seesaw bridge.
2 Calculations of ${\cal J}_{\nu}$
Taking account of the Euler-like block parametrization of the $6\times 6$ unitary flavor mixing matrix in the canonical seesaw mechanism [12], we have the exact expressions of $A$ and $R$ as follows:
$A=$ $\left(\begin{matrix}c_{14}c_{15}c_{16}&0&0\cr\vspace{-0.45cm}\cr\begin{array}[]{l}-c_{14}c_{15}\hat{s}_{16}\hat{s}^{*}_{26}-c_{14}\hat{s}_{15}\hat{s}^{*}_{25}c_{26}\\ -\hat{s}_{14}\hat{s}^{*}_{24}c_{25}c_{26}\end{array}&c_{24}c_{25}c_{26}&0\cr\vspace{-0.45cm}\cr\begin{array}[]{l}-c_{14}c_{15}\hat{s}_{16}c_{26}\hat{s}^{*}_{36}+c_{14}\hat{s}_{15}\hat{s}^{*}_{25}\hat{s}_{26}\hat{s}^{*}_{36}\\ -c_{14}\hat{s}_{15}c_{25}\hat{s}^{*}_{35}c_{36}+\hat{s}_{14}\hat{s}^{*}_{24}c_{25}\hat{s}_{26}\hat{s}^{*}_{36}\\ +\hat{s}_{14}\hat{s}^{*}_{24}\hat{s}_{25}\hat{s}^{*}_{35}c_{36}-\hat{s}_{14}c_{24}\hat{s}^{*}_{34}c_{35}c_{36}\end{array}&\begin{array}[]{l}\hskip 1.13791pt-c_{24}c_{25}\hat{s}_{26}\hat{s}^{*}_{36}-c_{24}\hat{s}_{25}\hat{s}^{*}_{35}c_{36}\\ \hskip 1.13791pt-\hat{s}_{24}\hat{s}^{*}_{34}c_{35}c_{36}\end{array}&c_{34}c_{35}c_{36}\cr\end{matrix}\right)\;,$ (4)
and
$R=$ $\left(\begin{matrix}\hat{s}^{*}_{14}c_{15}c_{16}&\hat{s}^{*}_{15}c_{16}&\hat{s}^{*}_{16}\cr\vspace{-0.45cm}\cr\begin{array}[]{l}-\hat{s}^{*}_{14}c_{15}\hat{s}_{16}\hat{s}^{*}_{26}-\hat{s}^{*}_{14}\hat{s}_{15}\hat{s}^{*}_{25}c_{26}\\ +c_{14}\hat{s}^{*}_{24}c_{25}c_{26}\end{array}&-\hat{s}^{*}_{15}\hat{s}_{16}\hat{s}^{*}_{26}+c_{15}\hat{s}^{*}_{25}c_{26}&c_{16}\hat{s}^{*}_{26}\cr\vspace{-0.45cm}\cr\begin{array}[]{l}-\hat{s}^{*}_{14}c_{15}\hat{s}_{16}c_{26}\hat{s}^{*}_{36}+\hat{s}^{*}_{14}\hat{s}_{15}\hat{s}^{*}_{25}\hat{s}_{26}\hat{s}^{*}_{36}\\ -\hat{s}^{*}_{14}\hat{s}_{15}c_{25}\hat{s}^{*}_{35}c_{36}-c_{14}\hat{s}^{*}_{24}c_{25}\hat{s}_{26}\hat{s}^{*}_{36}\\ -c_{14}\hat{s}^{*}_{24}\hat{s}_{25}\hat{s}^{*}_{35}c_{36}+c_{14}c_{24}\hat{s}^{*}_{34}c_{35}c_{36}\end{array}&\begin{array}[]{l}-\hat{s}^{*}_{15}\hat{s}_{16}c_{26}\hat{s}^{*}_{36}-c_{15}\hat{s}^{*}_{25}\hat{s}_{26}\hat{s}^{*}_{36}\\ +c_{15}c_{25}\hat{s}^{*}_{35}c_{36}\end{array}&c_{16}c_{26}\hat{s}^{*}_{36}\cr\end{matrix}\right)\;,\hskip 14.22636pt$ (5)
where $c_{ij}\equiv\cos\theta_{ij}$, $s_{ij}\equiv\sin\theta_{ij}$ and $\hat{s}_{ij}\equiv s_{ij}e^{{\rm i}\delta_{ij}}$ with $\theta_{ij}$ lying in the first quadrant. Some constraints on the 9 active-sterile flavor mixing angles $\theta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) can be obtained from the currently available bounds on non-unitarity of the PMNS lepton flavor mixing matrix $U$ [22, 23, 24, 25] via the relations $AA^{\dagger}=UU^{\dagger}$ and $RR^{\dagger}=I-AA^{\dagger}$. A conservative expectation is that the magnitudes of $\theta_{ij}$ are at least smaller than ${\cal O}(10^{-1.5})$, assuring that $A\simeq I$ and $U\simeq U_{0}$ hold to a good degree of accuracy, together with
$A^{-1}R\simeq R\simeq\left(\begin{matrix}\hat{s}^{*}_{14}&\hat{s}^{*}_{15}&\hat{s}^{*}_{16}\cr\hat{s}^{*}_{24}&\hat{s}^{*}_{25}&\hat{s}^{*}_{26}\cr\hat{s}^{*}_{34}&\hat{s}^{*}_{35}&\hat{s}^{*}_{36}\cr\end{matrix}\right)+{\cal O}(s^{3}_{ij})\;,$ (6)
where $s_{ij}\equiv\sin\theta_{ij}$ and $\hat{s}_{ij}\equiv s_{ij}e^{{\rm i}\delta_{ij}}$ with $\theta_{ij}$ lying in the first quadrant. A proper redefinition of the phases of three charged lepton fields in Eq. (1) allows one to rotate away 3 of the 9 phase parameters (or their independent combinations) of $A$ and $R$. We are therefore left with 6 nontrivial CP-violating phases which can be chosen from
$\alpha_{i}\equiv\delta_{i4}-\delta_{i5}\;,\quad\beta_{i}\equiv\delta_{i5}-\delta_{i6}\;,\quad\gamma_{i}\equiv\delta_{i6}-\delta_{i4}\;,$ (7)
as they satisfy $\alpha_{i}+\beta_{i}+\gamma_{i}=0$ (for $i=1,2,3$). Substituting Eq. (6) into Eq. (2) will make it possible for us to work out all the flavor parameters of the three light Majorana neutrinos. But in this work we concentrate on leptonic CP violation — a burning issue of particle physics which will definitely be established in the next-generation long-baseline neutrino oscillation experiments [29, 30].
We find that the most convenient way to calculate the Jarlskog invariant ${\cal J}_{\nu}$ defined in Eq. (3) is
${\cal J}_{\nu}=\frac{{\rm Im}\left[\left(M_{\nu}M^{\dagger}_{\nu}\right)_{e\mu}\left(M_{\nu}M^{\dagger}_{\nu}\right)_{\mu\tau}\left(M_{\nu}M^{\dagger}_{\nu}\right)_{\tau e}\right]}{\Delta_{21}\Delta_{31}\Delta_{32}}\;,$ (8)
where $M_{\nu}\equiv U_{0}D_{\nu}U^{T}_{0}$ and $\Delta_{ii^{\prime}}\equiv m^{2}_{i}-m^{2}_{i^{\prime}}$ (for $i,i^{\prime}=1,2,3$). One may derive the expressions of $m^{2}_{i}$ by use of the parameters of $D_{N}$ and $A^{-1}R$ from Eq. (2), and then arrive at those of $\Delta_{ii^{\prime}}$. But a global analysis of the neutrino oscillation data has given the best-fit values $\Delta_{21}\simeq 7.41\times 10^{-5}~{}{\rm eV}^{2}$ and $\Delta_{31}\simeq 2.51\times 10^{-3}~{}{\rm eV}^{2}$ (normal mass ordering) or $\Delta_{31}\simeq-2.41\times 10^{-3}~{}{\rm eV}^{2}$ (inverted mass ordering) [31, 32]. So the product $\Delta_{21}\Delta_{31}\Delta_{32}$ on the right-hand side of Eq. (8) is already known to us, and it is insensitive to the sign ambiguities of $\Delta_{31}$ and $\Delta_{32}$. In this case what we need to do is to calculate $(M_{\nu}M^{\dagger}_{\nu})_{e\mu}$, $(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}$ and $(M_{\nu}M^{\dagger}_{\nu})_{\tau e}$ in terms of the original seesaw parameters $M_{j}$, $\theta_{ij}$ and $\delta_{ij}$ (for $i=1,2,3$ and $j=4,5,6$) with the help of Eqs. (2) and (6):
$\left(M_{\nu}M^{\dagger}_{\nu}\right)_{e\mu}=$ $\hskip 2.84544pt\sum^{6}_{j=4}\sum^{6}_{k=4}M_{j}M_{k}I_{jk}\hat{s}^{*}_{1j}\hat{s}_{2k}\;,$
$\left(M_{\nu}M^{\dagger}_{\nu}\right)_{\mu\tau}=$ $\hskip 2.84544pt\sum^{6}_{j=4}\sum^{6}_{k=4}M_{j}M_{k}I_{jk}\hat{s}^{*}_{2j}\hat{s}_{3k}\;,$
$\left(M_{\nu}M^{\dagger}_{\nu}\right)_{\tau e}=$ $\hskip 2.84544pt\sum^{6}_{j=4}\sum^{6}_{k=4}M_{j}M_{k}I_{jk}\hat{s}^{*}_{3j}\hat{s}_{1k}\;,$ (9)
where
$I_{jk}\equiv\sum^{3}_{i=1}\hat{s}^{*}_{ij}\hat{s}_{ik}=I^{*}_{kj}\;,\;\left(j,k=4,5,6\right)\;.$ (10)
It is obvious that $I_{44}$, $I_{55}$ and $I_{66}$ are real and positive, while $I_{45}$, $I_{56}$ and $I_{64}$ depend respectively on the phase parameters $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$ (for $i=1,2,3$). Note that $I_{56}$ or $I_{64}$ can easily be obtained from $I_{45}$ by making the subscript replacements $(4,5)\to(5,6)$ or $(4,5)\to(6,4)$, together with $\alpha_{i}\to\beta_{i}$ or $\alpha_{i}\to\gamma_{i}$. Substituting Eq. (9) into Eq. (8) will then allow us to arrive at the explicit dependence of ${\cal J}_{\nu}$ on the 6 original CP-violating phases in the seesaw framework.
A very lengthy calculation shows that the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{6}_{j}$ (for $j=4,5,6$) and $M^{5}_{j}M_{k}$ (for $k\neq j$) do not contribute to ${\cal J}_{\nu}$ at all. But the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{3}_{j}M^{3}_{k}$ and $M^{4}_{j}M^{2}_{k}$ (for $k\neq j$) contribute to ${\cal J}_{\nu}$, so do the terms proportional to $M^{4}_{j}M_{k}M_{l}$, $M^{3}_{j}M^{2}_{k}M_{l}$ and $M^{2}_{j}M^{2}_{k}M^{2}_{l}$ (for $l\neq k\neq j$). So we proceed to present the analytical results of those nontrivial terms.
(1) For the sake of simplicity, we denote a sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{3}_{j}M^{3}_{k}$ (for $k\neq j$) as $T_{\bf 33}$, and a sum of those terms proportional to $M^{4}_{j}M^{2}_{k}$ (for $k\neq j$) as $T_{\bf 42}$. Our results of $T_{\bf 33}$ and $T_{\bf 42}$ are summarized as
$T_{\bf 33}=$ $\hskip 2.84544ptM^{3}_{4}M^{3}_{5}\left(I_{44}I_{55}-\left|I_{45}\right|^{2}\right)\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}4}s_{i^{\prime}5}\left(s^{2}_{i4}s^{2}_{i^{\prime\prime}5}+s^{2}_{i^{\prime\prime}4}s^{2}_{i^{\prime}5}-s^{2}_{i^{\prime}4}s^{2}_{i^{\prime\prime}5}-s^{2}_{i^{\prime\prime}4}s^{2}_{i5}\right)\sin\left(\alpha_{i}+\alpha_{i^{\prime}}\right)$
$-\hskip 2.27626pt\sum^{3}_{i=1}s^{2}_{i4}s^{2}_{i5}\left(s^{2}_{i^{\prime}4}s^{2}_{i^{\prime\prime}5}-s^{2}_{i^{\prime\prime}4}s^{2}_{i^{\prime}5}\right)\sin 2\alpha_{i}\Bigg{]}$
[…]
$T_{\bf 42}=$ $\hskip 2.84544ptM^{2}_{4}M^{2}_{5}\left(I_{44}I_{55}-\left|I_{45}\right|^{2}\right)\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}4}s_{i^{\prime}5}\left(M^{2}_{4}I_{44}s^{2}_{i^{\prime\prime}4}-M^{2}_{5}I_{55}s^{2}_{i^{\prime\prime}5}\right)\sin\left(\alpha_{i}-\alpha_{i^{\prime}}\right)\Bigg{]}$
[…]
(2) The terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ proportional to $M^{4}_{j}M_{k}M_{l}$ (for $l\neq k\neq j$) are summarily denoted as $T_{\bf 411}$. The explicit analytical result of $T_{\bf 411}$ is found to be
$T_{\bf 411}=$ $\hskip 3.69899ptM^{4}_{4}M_{5}M_{6}I_{44}s_{14}s_{24}s_{34}\Bigg{[}-s_{14}s_{24}s_{34}\sum^{3}_{i=1}s^{2}_{i5}\Big{[}s^{2}_{i^{\prime}6}\sin 2\left(\alpha_{i}+\gamma_{i^{\prime}}\right)-s^{2}_{i^{\prime\prime}6}\sin 2\left(\alpha_{i}+\gamma_{i^{\prime\prime}}\right)\Big{]}$
$+\sum^{3}_{i=1}s_{i4}\left(s^{2}_{i^{\prime}4}-s^{2}_{i^{\prime\prime}4}\right)\Big{[}s^{2}_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\sin\left(2\alpha_{i}+\gamma_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s^{2}_{i6}s_{i^{\prime}5}s_{i^{\prime\prime}5}\sin\left(\alpha_{i^{\prime}}+\alpha_{i^{\prime\prime}}+2\gamma_{i}\right)\Big{]}$
$+\sum^{3}_{i=1}s_{i4}s_{i^{\prime}5}s_{i^{\prime}6}\left(s^{2}_{i4}+s^{2}_{i^{\prime\prime}4}\right)\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}-\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}6}s_{i^{\prime\prime}5}\sin\left(\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}+\gamma_{i^{\prime}}\right)\Big{]}$
$+\sum^{3}_{i=1}s_{i4}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\left(s^{2}_{i4}+s^{2}_{i^{\prime}4}\right)\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}6}s_{i^{\prime\prime}5}\sin\left(\alpha_{i^{\prime\prime}}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}$
$+\sum^{3}_{i=1}2s_{i4}s_{i5}s_{i6}\left(s^{2}_{i^{\prime}4}+s^{2}_{i^{\prime\prime}4}\right)\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}-\beta_{i}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}6}s_{i^{\prime\prime}5}\sin\left(\alpha_{i^{\prime\prime}}-\beta_{i}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$+{\rm term}\big{\{}\left(4,5,6\right)\to\left(5,4,6\right);\;\left(\alpha_{i},\;\beta_{i},\;\gamma_{i}\right)\to-\left(\alpha_{i},\;\gamma_{i},\;\beta_{i}\right)\big{\}}$
$+{\rm term}\big{\{}\left(4,5,6\right)\to\left(6,5,4\right);\;\left(\alpha_{i},\;\beta_{i},\;\gamma_{i}\right)\to-\left(\beta_{i},\;\alpha_{i},\;\gamma_{i}\right)\big{\}}\;,$ (13)
where the three light subscripts $\left(i,i^{\prime},i^{\prime\prime}\right)$ run cyclically over $\left(1,2,3\right)$, and the heavy subscript replacements for those abbreviation terms have the same implications as in Eqs. (11) and (12).
(3) A sum of the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{3}_{4}M^{2}_{5}M_{6}$, $M^{3}_{4}M^{2}_{6}M_{5}$, $M^{3}_{5}M^{2}_{4}M_{6}$, $M^{3}_{5}M^{2}_{6}M_{4}$, $M^{3}_{6}M^{2}_{4}M_{5}$ and $M^{3}_{6}M^{2}_{5}M_{4}$ is denoted as $T_{\bf 321}$. The explicit analytical result of $T_{\bf 321}$ is obtained as follows:
$T_{\bf 321}=$ $\hskip 3.69899ptM^{3}_{4}M^{2}_{5}M_{6}\Bigg{\{}\left(I_{44}I_{55}-|I_{45}|^{2}\right)\Bigg{[}\sum^{3}_{i=1}s^{2}_{i4}s_{i^{\prime}5}s_{i^{\prime\prime}5}\Big{[}s_{i4}s_{i6}s_{i^{\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}+\beta_{i^{\prime\prime}}-\gamma_{i}\right)-s_{i4}s_{i6}s_{i^{\prime}6}s_{i^{\prime\prime}4}$
×$\times\sin\left(\alpha_{i^{\prime\prime}}+\beta_{i^{\prime}}-\gamma_{i}\right)+s^{2}_{i^{\prime}4}s_{i^{\prime}6}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}+\beta_{i^{\prime\prime}}-\gamma_{i^{\prime}}\right)-s_{i^{\prime}4}s^{2}_{i^{\prime}6}s_{i^{\prime\prime}4}\sin\left(\alpha_{i^{\prime\prime}}+\beta_{i^{\prime}}-\gamma_{i^{\prime}}\right)$
$+s_{i^{\prime}4}s_{i^{\prime\prime}4}s^{2}_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}+\beta_{i^{\prime\prime}}-\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}6}s^{2}_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime\prime}}+\beta_{i^{\prime}}-\gamma_{i^{\prime\prime}}\right)\Big{]}\Bigg{]}$
$+I_{44}s_{14}s_{24}s_{34}\Bigg{[}2\sum^{3}_{i=1}s_{i6}s_{i^{\prime}5}s_{i^{\prime\prime}5}\Big{[}s_{i4}s_{i5}s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\alpha_{i}+\beta_{i^{\prime}}-\gamma_{i}\right)+s_{i4}s_{i5}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}$
×$\times\cos\left(\alpha_{i}+\beta_{i^{\prime\prime}}-\gamma_{i}\right)+s_{i^{\prime}4}s_{i^{\prime}5}s_{i5}s_{i6}\cos\left(\alpha_{i^{\prime}}+\beta_{i}-\gamma_{i}\right)+s_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\cos\left(\alpha_{i^{\prime}}+\beta_{i^{\prime\prime}}-\gamma_{i}\right)$
$+s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i5}s_{i6}\cos\left(\alpha_{i^{\prime\prime}}+\beta_{i}-\gamma_{i}\right)+s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\alpha_{i^{\prime\prime}}+\beta_{i^{\prime}}-\gamma_{i}\right)$
$-s_{i4}s_{i6}\left(s^{2}_{i^{\prime}5}+s^{2}_{i^{\prime\prime}5}\right)\cos 2\gamma_{i}-s_{i^{\prime}4}s_{i^{\prime}6}\left(s^{2}_{i5}+s^{2}_{i^{\prime\prime}5}\right)\cos\left(\gamma_{i}+\gamma_{i^{\prime}}\right)$
$-s_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\left(s^{2}_{i5}+s^{2}_{i^{\prime}5}\right)\cos\left(\gamma_{i}+\gamma_{i^{\prime\prime}}\right)\Big{]}\sin\left(\alpha_{i^{\prime\prime}}-\alpha_{i^{\prime}}\right)\Bigg{]}$
$+I_{44}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i6}\left(s^{2}_{i^{\prime}5}s^{2}_{i^{\prime\prime}4}-s^{2}_{i^{\prime}4}s^{2}_{i^{\prime\prime}5}\right)\Big{[}s_{i4}s_{i5}s_{i^{\prime}5}s_{i^{\prime}6}\sin\left(\alpha_{i}+\beta_{i^{\prime}}-\gamma_{i}\right)+s_{i4}s_{i5}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}$
×$\times\sin\left(\alpha_{i}+\beta_{i^{\prime\prime}}-\gamma_{i}\right)+s_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}+\beta_{i^{\prime\prime}}-\gamma_{i}\right)+s_{i^{\prime}4}s_{i^{\prime}5}s_{i5}s_{i6}\sin\left(\alpha_{i^{\prime}}+\beta_{i}-\gamma_{i}\right)$
$+s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i5}s_{i6}\sin\left(\alpha_{i^{\prime\prime}}+\beta_{i}-\gamma_{i}\right)+s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i^{\prime}5}s_{i^{\prime}6}\sin\left(\alpha_{i^{\prime\prime}}+\beta_{i^{\prime}}-\gamma_{i}\right)$
$-s_{i4}s_{i6}\left(s^{2}_{i5}-I^{-1}_{44}\left|I_{45}\right|^{2}\right)\sin 2\gamma_{i}-s_{i^{\prime}4}s_{i^{\prime}6}\left(s^{2}_{i^{\prime}5}-I^{-1}_{44}\left|I_{45}\right|^{2}\right)\sin\left(\gamma_{i}+\gamma_{i^{\prime}}\right)$
$-s_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\left(s^{2}_{i^{\prime\prime}5}-I^{-1}_{44}\left|I_{45}\right|^{2}\right)\sin\left(\gamma_{i}+\gamma_{i^{\prime\prime}}\right)\Big{]}\Bigg{]}$
$+\Bigg{[}\sum^{3}_{i=1}s_{i5}\Big{[}s_{i6}\left(s^{2}_{i^{\prime}4}+s^{2}_{i^{\prime\prime}4}\right)\cos\beta_{i}-s_{i4}s_{i^{\prime}4}s_{i^{\prime}6}\cos\left(\alpha_{i}+\gamma_{i^{\prime}}\right)-s_{i4}s_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\cos\left(\alpha_{i}+\gamma_{i^{\prime\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s^{2}_{i4}s_{i^{\prime}5}s_{i^{\prime\prime}5}\Big{[}s_{i4}s_{i5}s_{i^{\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i}+\alpha_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)-s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}4}\sin\left(\alpha_{i}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)$
$+s^{2}_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime\prime}6}\sin\left(2\alpha_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime}6}s_{i^{\prime\prime}4}\sin\left(\alpha_{i^{\prime}}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)$
$+s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i^{\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i^{\prime}}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime\prime}}\right)-s_{i^{\prime}6}s^{2}_{i^{\prime\prime}4}s_{i^{\prime\prime}5}\sin\left(2\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$+\Bigg{[}\sum^{3}_{i=1}s^{2}_{i4}s_{i^{\prime}5}s_{i^{\prime\prime}5}\Big{[}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}4}\cos\left(\alpha_{i}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)-s_{i4}s_{i5}s_{i^{\prime}4}s_{i^{\prime\prime}6}\cos\left(\alpha_{i}+\alpha_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)$
$+s_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime}6}s_{i^{\prime\prime}4}\cos\left(\alpha_{i^{\prime}}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)-s^{2}_{i^{\prime}4}s_{i^{\prime}5}s_{i^{\prime\prime}6}\cos\left(2\alpha_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)$
$+s_{i^{\prime}6}s^{2}_{i^{\prime\prime}4}s_{i^{\prime\prime}5}\cos\left(2\alpha_{i^{\prime\prime}}-\beta_{i^{\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}4}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\cos\left(\alpha_{i^{\prime}}+\alpha_{i^{\prime\prime}}-\beta_{i^{\prime\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i5}\Big{[}s_{i6}\left(s^{2}_{i^{\prime}4}+s^{2}_{i^{\prime\prime}4}\right)\sin\beta_{i}+s_{i4}s_{i^{\prime}4}s_{i^{\prime}6}\sin\left(\alpha_{i}+\gamma_{i^{\prime}}\right)+s_{i4}s_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i}+\gamma_{i^{\prime\prime}}\right)\Big{]}\Bigg{]}\Bigg{\}}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(4,6,5\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to-\left(\gamma_{i},\beta_{i},\alpha_{i}\right)\right\}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(5,4,6\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to-\left(\alpha_{i},\gamma_{i},\beta_{i}\right)\right\}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(5,6,4\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to+\left(\beta_{i},\gamma_{i},\alpha_{i}\right)\right\}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(6,4,5\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to+\left(\gamma_{i},\alpha_{i},\beta_{i}\right)\right\}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(6,5,4\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to-\left(\beta_{i},\alpha_{i},\gamma_{i}\right)\right\}\;,$ (14)
where the subscripts $\left(i,i^{\prime},i^{\prime\prime}\right)$ run cyclically over $\left(1,2,3\right)$, and the heavy subscript replacements for those abbreviation terms have the same implications as in Eq. (13).
(4) Finally, the terms of ${\rm Im}\left[(M_{\nu}M^{\dagger}_{\nu})_{e\mu}(M_{\nu}M^{\dagger}_{\nu})_{\mu\tau}(M_{\nu}M^{\dagger}_{\nu})_{\tau e}\right]$ that are proportional to $M^{2}_{4}M^{2}_{5}M^{2}_{6}$ are summarily denoted as $T_{\bf 222}$. The explicit analytical result of $T_{\bf 222}$ is found to be
$T_{\bf 222}=$ $\hskip 3.69899ptM^{2}_{4}M^{2}_{5}M^{2}_{6}\Bigg{\{}I_{44}I_{55}I_{66}s_{14}s_{15}s_{26}s_{36}\Big{[}s_{25}s_{34}\sin\left(\alpha_{1}+\beta_{2}+\gamma_{3}\right)-s_{24}s_{35}\sin\left(\alpha_{1}+\beta_{3}+\gamma_{2}\right)\Big{]}$
$+\left(\left|I_{45}\right|^{2}I_{66}-\left|I_{64}\right|^{2}I_{55}\right)\Bigg{[}\sum^{3}_{i=1}s^{2}_{i4}s_{i^{\prime}5}s_{i^{\prime}6}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\sin\left(\beta_{i^{\prime}}-\beta_{i^{\prime\prime}}\right)\Bigg{]}$
$+I_{44}\Bigg{[}\sum^{3}_{i=1}s_{i5}s_{i6}\Big{[}s_{i5}s_{i6}\cos 2\beta_{i}+2s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\beta_{i}+\beta_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\sin\left(\alpha_{i}-\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$-I_{44}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s_{i^{\prime}6}s_{i^{\prime\prime}6}\Big{[}s_{i^{\prime}5}s_{i^{\prime\prime}4}\cos\left(\alpha_{i}-\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)-s_{i^{\prime}4}s_{i^{\prime\prime}5}\sin\left(\alpha_{i}-\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i5}s_{i6}\Big{[}s_{i5}s_{i6}\sin 2\beta_{i}+2s_{i^{\prime}5}s_{i^{\prime}6}\sin\left(\beta_{i}+\beta_{i^{\prime}}\right)\Big{]}\Bigg{]}$
$+2s_{14}s_{15}s_{26}s_{36}\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}s^{2}_{i6}\cos\left(\alpha_{1}+\alpha_{i}\right)+\sum^{3}_{i=1}s_{i6}s_{i^{\prime}6}\Big{[}s_{i4}s_{i^{\prime}5}\cos\left(\alpha_{1}-\beta_{i^{\prime}}-\gamma_{i}\right)$
$+s_{i5}s_{i^{\prime}4}\cos\left(\alpha_{1}-\beta_{i}-\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}s_{25}s_{34}\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{i}-\beta_{2}-\gamma_{3}\right)-s_{24}s_{35}\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{i}-\beta_{3}-\gamma_{2}\right)\Bigg{]}$
$-2s_{14}s_{15}s_{26}s_{36}\Bigg{[}s_{25}s_{34}\sum^{3}_{i=1}s_{i4}s_{i6}\Big{[}s_{i5}s_{i6}\cos\left(\beta_{i}+\beta_{2}-\gamma_{i}-\gamma_{3}\right)+s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\beta_{i^{\prime}}+\beta_{2}-\gamma_{i}-\gamma_{3}\right)$
$+s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\cos\left(\beta_{i^{\prime\prime}}+\beta_{2}-\gamma_{i}-\gamma_{3}\right)\Big{]}-s_{24}s_{35}\sum^{3}_{i=1}s_{i4}s_{i6}\Big{[}s_{i5}s_{i6}\cos\left(\beta_{i}+\beta_{3}-\gamma_{i}-\gamma_{2}\right)$
$+s_{i^{\prime}5}s_{i^{\prime}6}\cos\left(\beta_{i^{\prime}}+\beta_{3}-\gamma_{i}-\gamma_{2}\right)+s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\cos\left(\beta_{i^{\prime\prime}}+\beta_{3}-\gamma_{i}-\gamma_{2}\right)\Big{]}\Bigg{]}$
×$\times\Bigg{[}\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{1}+\alpha_{i}\right)\Bigg{]}$
$+\left(s^{2}_{14}s^{2}_{25}-s^{2}_{15}s^{2}_{24}\right)s^{2}_{36}\Bigg{[}\sum^{3}_{i=1}\Big{[}s_{i4}s_{i5}s_{i^{\prime}4}s_{i^{\prime}5}\left(s^{2}_{i6}-s^{2}_{i^{\prime}6}\right)\sin\left(\alpha_{i}-\alpha_{i^{\prime}}\right)+s_{i5}s_{i6}s_{i^{\prime}5}s_{i^{\prime}6}\left(s^{2}_{i4}-s^{2}_{i^{\prime}4}\right)$
×$\times\sin\left(\beta_{i}-\beta_{i^{\prime}}\right)+s_{i4}s_{i6}s_{i^{\prime}4}s_{i^{\prime}6}\left(s^{2}_{i5}-s^{2}_{i^{\prime}5}\right)\sin\left(\gamma_{i}-\gamma_{i^{\prime}}\right)\Big{]}$
$-\sum^{3}_{i=1}s_{i4}s_{i5}\Big{[}s_{i^{\prime}5}s_{i^{\prime}6}s_{i^{\prime\prime}4}s_{i^{\prime\prime}6}\sin\left(\alpha_{i}+\beta_{i^{\prime}}+\gamma_{i^{\prime\prime}}\right)+s_{i^{\prime}4}s_{i^{\prime}6}s_{i^{\prime\prime}5}s_{i^{\prime\prime}6}\sin\left(\alpha_{i}+\beta_{i^{\prime\prime}}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}\Bigg{\}}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(4,6,5\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to-\left(\gamma_{i},\beta_{i},\alpha_{i}\right)\right\}$
$+{\rm term}\left\{\left(4,5,6\right)\to\left(6,5,4\right);\;\left(\alpha_{i},\beta_{i},\gamma_{i}\right)\to-\left(\beta_{i},\alpha_{i},\gamma_{i}\right)\right\},$ (15)
where the three light subscripts $\left(i,i^{\prime},i^{\prime\prime}\right)$ run cyclically over $\left(1,2,3\right)$, and the heavy subscript replacements for those abbreviation terms have the same implications as in Eq. (14). It is straightforward to see that $T_{\bf 411}$, $T_{\bf 321}$ and $T_{\bf 222}$ characterize the three-species interference effects in which all the three kinds of heavy Majorana neutrinos take part, and hence their analytical expressions are much more complicated than those of $T_{\bf 33}$ and $T_{\bf 42}$.
To summarize, a general and explicit expression of the Jarlskog invariant ${\cal J}_{\nu}$ in the canonical seesaw mechanism turns out to be
${\cal J}_{\nu}=\frac{T_{\bf 33}+T_{\bf 42}+T_{\bf 411}+T_{\bf 321}+T_{\bf 222}}{\Delta_{21}\Delta_{31}\Delta_{32}}\;,$ (16)
where the analytical results of $T_{\bf 33}$, $T_{\bf 42}$, $T_{\bf 411}$, $T_{\bf 321}$ and $T_{\bf 222}$ have been given in Eqs. (11)—(15). Let us emphasize that these results are completely new. It is certainly too lengthy to write out all the abbreviation terms in the above equations, but one may figure out all the combinations of the CP-violating phases $\alpha_{i}$, $\beta_{i}$ and $\gamma_{i}$ (for $i=1,2,3$) that appear in the expression of ${\cal J}_{\nu}$. After a very tedious survey of all the possibilities, we find that there are totally 240 linear combinations of the 6 original seesaw phase parameters:
$\begin{array}[]{lllllllll}\sin\alpha_{1},&\sin\alpha_{2},&\sin\alpha_{3};&\sin\beta_{1},&\sin\beta_{2},&\sin\beta_{3};&\sin\gamma_{1},&\sin\gamma_{2},&\sin\gamma_{3};\\ \vspace{-0.4cm}\\ \sin 2\alpha_{1},&\sin 2\alpha_{2},&\sin 2\alpha_{3};&\sin 2\beta_{1},&\sin 2\beta_{2},&\sin 2\beta_{3};&\sin 2\gamma_{1},&\sin 2\gamma_{2},&\sin 2\gamma_{3};\end{array}$
$\begin{array}[]{llllll}\sin\left(\alpha_{1}+\alpha_{2}\right),&\sin\left(\alpha_{2}+\alpha_{3}\right),&\sin\left(\alpha_{3}+\alpha_{1}\right);&\sin\left(\beta_{1}+\beta_{2}\right),&\sin\left(\beta_{2}+\beta_{3}\right),&\sin\left(\beta_{3}+\beta_{1}\right);\\ \sin\left(\gamma_{1}+\gamma_{2}\right),&\sin\left(\gamma_{2}+\gamma_{3}\right),&\sin\left(\gamma_{3}+\gamma_{1}\right);&&&\\ \sin\left(\alpha_{1}-\alpha_{2}\right),&\sin\left(\alpha_{2}-\alpha_{3}\right),&\sin\left(\alpha_{3}-\alpha_{1}\right);&\sin\left(\beta_{1}-\beta_{2}\right),&\sin\left(\beta_{2}-\beta_{3}\right),&\sin\left(\beta_{3}-\beta_{1}\right);\\ \sin\left(\gamma_{1}-\gamma_{2}\right),&\sin\left(\gamma_{2}-\gamma_{3}\right),&\sin\left(\gamma_{3}-\gamma_{1}\right);&&&\\ \vspace{-0.3cm}\\ \sin\left(\alpha_{1}+\beta_{2}\right),&\sin\left(\alpha_{1}+\beta_{3}\right),&\sin\left(\alpha_{2}+\beta_{1}\right),&\sin\left(\alpha_{2}+\beta_{3}\right),&\sin\left(\alpha_{3}+\beta_{1}\right),&\sin\left(\alpha_{3}+\beta_{2}\right);\\ \sin\left(\alpha_{1}+\gamma_{2}\right),&\sin\left(\alpha_{1}+\gamma_{3}\right),&\sin\left(\alpha_{2}+\gamma_{1}\right),&\sin\left(\alpha_{2}+\gamma_{3}\right),&\sin\left(\alpha_{3}+\gamma_{1}\right),&\sin\left(\alpha_{3}+\gamma_{2}\right);\\ \sin\left(\beta_{1}+\gamma_{2}\right),&\sin\left(\beta_{1}+\gamma_{3}\right),&\sin\left(\beta_{2}+\gamma_{1}\right),&\sin\left(\beta_{2}+\gamma_{3}\right),&\sin\left(\beta_{3}+\gamma_{1}\right),&\sin\left(\beta_{3}+\gamma_{2}\right);\\ \vspace{-0.3cm}\\ \sin 2\left(\alpha_{1}+\beta_{2}\right),&\sin 2\left(\alpha_{1}+\beta_{3}\right),&\sin 2\left(\alpha_{2}+\beta_{1}\right),&\sin 2\left(\alpha_{2}+\beta_{3}\right),&\sin 2\left(\alpha_{3}+\beta_{1}\right),&\sin 2\left(\alpha_{3}+\beta_{2}\right);\\ \sin 2\left(\alpha_{1}+\gamma_{2}\right),&\sin 2\left(\alpha_{1}+\gamma_{3}\right),&\sin 2\left(\alpha_{2}+\gamma_{1}\right),&\sin 2\left(\alpha_{2}+\gamma_{3}\right),&\sin 2\left(\alpha_{3}+\gamma_{1}\right),&\sin 2\left(\alpha_{3}+\gamma_{2}\right);\\ \sin 2\left(\beta_{1}+\gamma_{2}\right),&\sin 2\left(\beta_{1}+\gamma_{3}\right),&\sin 2\left(\beta_{2}+\gamma_{1}\right),&\sin 2\left(\beta_{2}+\gamma_{3}\right),&\sin 2\left(\beta_{3}+\gamma_{1}\right),&\sin 2\left(\beta_{3}+\gamma_{2}\right);\\ \vspace{-0.3cm}\\ \sin\left(2\alpha_{1}-\beta_{2}\right),&\sin\left(2\alpha_{1}-\beta_{3}\right),&\sin\left(2\alpha_{2}-\beta_{1}\right),&\sin\left(2\alpha_{2}-\beta_{3}\right),&\sin\left(2\alpha_{3}-\beta_{1}\right),&\sin\left(2\alpha_{3}-\beta_{2}\right);\\ \sin\left(2\alpha_{1}-\gamma_{2}\right),&\sin\left(2\alpha_{1}-\gamma_{3}\right),&\sin\left(2\alpha_{2}-\gamma_{1}\right),&\sin\left(2\alpha_{2}-\gamma_{3}\right),&\sin\left(2\alpha_{3}-\gamma_{1}\right),&\sin\left(2\alpha_{3}-\gamma_{2}\right);\\ \sin\left(2\beta_{1}-\alpha_{2}\right),&\sin\left(2\beta_{1}-\alpha_{3}\right),&\sin\left(2\beta_{2}-\alpha_{1}\right),&\sin\left(2\beta_{2}-\alpha_{3}\right),&\sin\left(2\beta_{3}-\alpha_{1}\right),&\sin\left(2\beta_{3}-\alpha_{2}\right);\\ \sin\left(2\beta_{1}-\gamma_{2}\right),&\sin\left(2\beta_{1}-\gamma_{3}\right),&\sin\left(2\beta_{2}-\gamma_{1}\right),&\sin\left(2\beta_{2}-\gamma_{3}\right),&\sin\left(2\beta_{3}-\gamma_{1}\right),&\sin\left(2\beta_{3}-\gamma_{2}\right);\\ \sin\left(2\gamma_{1}-\alpha_{2}\right),&\sin\left(2\gamma_{1}-\alpha_{3}\right),&\sin\left(2\gamma_{2}-\alpha_{1}\right),&\sin\left(2\gamma_{2}-\alpha_{3}\right),&\sin\left(2\gamma_{3}-\alpha_{1}\right),&\sin\left(2\gamma_{3}-\alpha_{2}\right);\\ \sin\left(2\gamma_{1}-\beta_{2}\right),&\sin\left(2\gamma_{1}-\beta_{3}\right),&\sin\left(2\gamma_{2}-\beta_{1}\right),&\sin\left(2\gamma_{2}-\beta_{3}\right),&\sin\left(2\gamma_{3}-\beta_{1}\right),&\sin\left(2\gamma_{3}-\beta_{2}\right);\end{array}$
$\begin{array}[]{llll}\sin\left(\alpha_{1}+\alpha_{2}+2\beta_{3}\right),&\sin\left(\alpha_{1}+\alpha_{2}+2\gamma_{3}\right),&\sin\left(\alpha_{1}+\alpha_{3}+2\beta_{2}\right),&\sin\left(\alpha_{1}+\alpha_{3}+2\gamma_{2}\right),\\ \sin\left(\alpha_{2}+\alpha_{3}+2\beta_{1}\right),&\sin\left(\alpha_{2}+\alpha_{3}+2\gamma_{1}\right);&\sin\left(\beta_{1}+\beta_{2}+2\alpha_{3}\right),&\sin\left(\beta_{1}+\beta_{2}+2\gamma_{3}\right),\\ \sin\left(\beta_{1}+\beta_{3}+2\alpha_{2}\right),&\sin\left(\beta_{1}+\beta_{3}+2\gamma_{2}\right),&\sin\left(\beta_{2}+\beta_{3}+2\alpha_{1}\right),&\sin\left(\beta_{2}+\beta_{3}+2\gamma_{1}\right);\\ \sin\left(\gamma_{1}+\gamma_{2}+2\alpha_{3}\right),&\sin\left(\gamma_{1}+\gamma_{2}+2\beta_{3}\right),&\sin\left(\gamma_{1}+\gamma_{3}+2\alpha_{2}\right),&\sin\left(\gamma_{1}+\gamma_{3}+2\beta_{2}\right),\\ \sin\left(\gamma_{2}+\gamma_{3}+2\alpha_{1}\right),&\sin\left(\gamma_{2}+\gamma_{3}+2\beta_{1}\right);&&\\ \vspace{-0.3cm}\\ \sin\left(\alpha_{1}+\beta_{2}+\gamma_{3}\right),&\sin\left(\alpha_{1}+\beta_{3}+\gamma_{2}\right),&\sin\left(\alpha_{2}+\beta_{1}+\gamma_{3}\right),&\sin\left(\alpha_{2}+\beta_{3}+\gamma_{1}\right),\\ \sin\left(\alpha_{3}+\beta_{1}+\gamma_{2}\right),&\sin\left(\alpha_{3}+\beta_{2}+\gamma_{1}\right).&&\end{array}$
$\begin{array}[]{llll}\sin\left(\alpha_{1}+\alpha_{2}-\beta_{1}\right),&\sin\left(\alpha_{1}+\alpha_{2}-\beta_{2}\right),&\sin\left(\alpha_{1}+\alpha_{2}-\beta_{3}\right),&\sin\left(\alpha_{1}+\alpha_{3}-\beta_{1}\right),\\ \sin\left(\alpha_{1}+\alpha_{3}-\beta_{2}\right),&\sin\left(\alpha_{1}+\alpha_{3}-\beta_{3}\right),&\sin\left(\alpha_{2}+\alpha_{3}-\beta_{1}\right),&\sin\left(\alpha_{2}+\alpha_{3}-\beta_{2}\right),\\ \sin\left(\alpha_{2}+\alpha_{3}-\beta_{3}\right);&\sin\left(\alpha_{1}+\alpha_{2}-\gamma_{1}\right),&\sin\left(\alpha_{1}+\alpha_{2}-\gamma_{2}\right),&\sin\left(\alpha_{1}+\alpha_{2}-\gamma_{3}\right),\\ \sin\left(\alpha_{1}+\alpha_{3}-\gamma_{1}\right),&\sin\left(\alpha_{1}+\alpha_{3}-\gamma_{2}\right),&\sin\left(\alpha_{1}+\alpha_{3}-\gamma_{3}\right),&\sin\left(\alpha_{2}+\alpha_{3}-\gamma_{1}\right),\\ \sin\left(\alpha_{2}+\alpha_{3}-\gamma_{2}\right),&\sin\left(\alpha_{2}+\alpha_{3}-\gamma_{3}\right);&\sin\left(\beta_{1}+\beta_{2}-\alpha_{1}\right),&\sin\left(\beta_{1}+\beta_{2}-\alpha_{2}\right),\\ \sin\left(\beta_{1}+\beta_{2}-\alpha_{3}\right),&\sin\left(\beta_{1}+\beta_{3}-\alpha_{1}\right),&\sin\left(\beta_{1}+\beta_{3}-\alpha_{2}\right),&\sin\left(\beta_{1}+\beta_{3}-\alpha_{3}\right),\\ \sin\left(\beta_{2}+\beta_{3}-\alpha_{1}\right),&\sin\left(\beta_{2}+\beta_{3}-\alpha_{2}\right),&\sin\left(\beta_{2}+\beta_{3}-\alpha_{3}\right);&\sin\left(\beta_{1}+\beta_{2}-\gamma_{1}\right),\\ \sin\left(\beta_{1}+\beta_{2}-\gamma_{2}\right),&\sin\left(\beta_{1}+\beta_{2}-\gamma_{3}\right),&\sin\left(\beta_{1}+\beta_{3}-\gamma_{1}\right),&\sin\left(\beta_{1}+\beta_{3}-\gamma_{2}\right),\\ \sin\left(\beta_{1}+\beta_{3}-\gamma_{3}\right),&\sin\left(\beta_{2}+\beta_{3}-\gamma_{1}\right),&\sin\left(\beta_{2}+\beta_{3}-\gamma_{2}\right),&\sin\left(\beta_{2}+\beta_{3}-\gamma_{3}\right);\\ \sin\left(\gamma_{1}+\gamma_{2}-\alpha_{1}\right),&\sin\left(\gamma_{1}+\gamma_{2}-\alpha_{2}\right),&\sin\left(\gamma_{1}+\gamma_{2}-\alpha_{3}\right),&\sin\left(\gamma_{1}+\gamma_{3}-\alpha_{1}\right),\\ \sin\left(\gamma_{1}+\gamma_{3}-\alpha_{2}\right),&\sin\left(\gamma_{1}+\gamma_{3}-\alpha_{3}\right),&\sin\left(\gamma_{2}+\gamma_{3}-\alpha_{1}\right),&\sin\left(\gamma_{2}+\gamma_{3}-\alpha_{2}\right),\\ \sin\left(\gamma_{2}+\gamma_{3}-\alpha_{3}\right);&\sin\left(\gamma_{1}+\gamma_{2}-\beta_{1}\right),&\sin\left(\gamma_{1}+\gamma_{2}-\beta_{2}\right),&\sin\left(\gamma_{1}+\gamma_{2}-\beta_{3}\right),\\ \sin\left(\gamma_{1}+\gamma_{3}-\beta_{1}\right),&\sin\left(\gamma_{1}+\gamma_{3}-\beta_{2}\right),&\sin\left(\gamma_{1}+\gamma_{3}-\beta_{3}\right),&\sin\left(\gamma_{2}+\gamma_{3}-\beta_{1}\right),\\ \sin\left(\gamma_{2}+\gamma_{3}-\beta_{2}\right),&\sin\left(\gamma_{2}+\gamma_{3}-\beta_{3}\right);&&\\ \vspace{-0.3cm}\\ \sin\left(\alpha_{1}+\beta_{2}-\gamma_{1}\right),&\sin\left(\alpha_{1}+\beta_{2}-\gamma_{2}\right),&\sin\left(\alpha_{1}+\beta_{2}-\gamma_{3}\right),&\sin\left(\alpha_{1}+\beta_{3}-\gamma_{1}\right),\\ \sin\left(\alpha_{1}+\beta_{3}-\gamma_{2}\right),&\sin\left(\alpha_{1}+\beta_{3}-\gamma_{3}\right),&\sin\left(\alpha_{2}+\beta_{1}-\gamma_{1}\right),&\sin\left(\alpha_{2}+\beta_{1}-\gamma_{2}\right),\\ \sin\left(\alpha_{2}+\beta_{1}-\gamma_{3}\right),&\sin\left(\alpha_{2}+\beta_{3}-\gamma_{1}\right),&\sin\left(\alpha_{2}+\beta_{3}-\gamma_{2}\right),&\sin\left(\alpha_{2}+\beta_{3}-\gamma_{3}\right),\\ \sin\left(\alpha_{3}+\beta_{1}-\gamma_{1}\right),&\sin\left(\alpha_{3}+\beta_{1}-\gamma_{2}\right),&\sin\left(\alpha_{3}+\beta_{1}-\gamma_{3}\right),&\sin\left(\alpha_{3}+\beta_{2}-\gamma_{1}\right),\\ \sin\left(\alpha_{3}+\beta_{2}-\gamma_{2}\right),&\sin\left(\alpha_{3}+\beta_{2}-\gamma_{3}\right);&\sin\left(\alpha_{1}-\beta_{1}+\gamma_{2}\right),&\sin\left(\alpha_{1}-\beta_{1}+\gamma_{3}\right),\\ \sin\left(\alpha_{1}-\beta_{2}+\gamma_{2}\right),&\sin\left(\alpha_{1}-\beta_{2}+\gamma_{3}\right),&\sin\left(\alpha_{1}-\beta_{3}+\gamma_{2}\right),&\sin\left(\alpha_{1}-\beta_{3}+\gamma_{3}\right),\\ \sin\left(\alpha_{2}-\beta_{1}+\gamma_{1}\right),&\sin\left(\alpha_{2}-\beta_{1}+\gamma_{3}\right),&\sin\left(\alpha_{2}-\beta_{2}+\gamma_{1}\right),&\sin\left(\alpha_{2}-\beta_{2}+\gamma_{3}\right),\\ \sin\left(\alpha_{2}-\beta_{3}+\gamma_{1}\right),&\sin\left(\alpha_{2}-\beta_{3}+\gamma_{3}\right),&\sin\left(\alpha_{3}-\beta_{1}+\gamma_{1}\right),&\sin\left(\alpha_{3}-\beta_{1}+\gamma_{2}\right),\\ \sin\left(\alpha_{3}-\beta_{2}+\gamma_{1}\right),&\sin\left(\alpha_{3}-\beta_{2}+\gamma_{2}\right),&\sin\left(\alpha_{3}-\beta_{3}+\gamma_{1}\right),&\sin\left(\alpha_{3}-\beta_{3}+\gamma_{2}\right);\\ \sin\left(\alpha_{1}-\beta_{1}-\gamma_{2}\right),&\sin\left(\alpha_{1}-\beta_{1}-\gamma_{3}\right),&\sin\left(\alpha_{1}-\beta_{2}-\gamma_{1}\right),&\sin\left(\alpha_{1}-\beta_{2}-\gamma_{3}\right),\\ \sin\left(\alpha_{1}-\beta_{3}-\gamma_{1}\right),&\sin\left(\alpha_{1}-\beta_{3}-\gamma_{2}\right),&\sin\left(\alpha_{2}-\beta_{1}-\gamma_{2}\right),&\sin\left(\alpha_{2}-\beta_{1}-\gamma_{3}\right),\\ \sin\left(\alpha_{2}-\beta_{2}-\gamma_{1}\right),&\sin\left(\alpha_{2}-\beta_{2}-\gamma_{3}\right),&\sin\left(\alpha_{2}-\beta_{3}-\gamma_{1}\right),&\sin\left(\alpha_{2}-\beta_{3}-\gamma_{2}\right),\\ \sin\left(\alpha_{3}-\beta_{1}-\gamma_{2}\right),&\sin\left(\alpha_{3}-\beta_{1}-\gamma_{3}\right),&\sin\left(\alpha_{3}-\beta_{2}-\gamma_{1}\right),&\sin\left(\alpha_{3}-\beta_{2}-\gamma_{3}\right),\\ \sin\left(\alpha_{3}-\beta_{3}-\gamma_{1}\right),&\sin\left(\alpha_{3}-\beta_{3}-\gamma_{2}\right).&&\end{array}$
Note that the number of such linear phase combinations will not be reduced if $\gamma_{i}=-\left(\alpha_{i}+\beta_{i}\right)$ is taken into account. That is why we have kept $\gamma_{1}$, $\gamma_{2}$ and $\gamma_{3}$ in the above list as in Eqs. (11)—(15), just for the sake of simplicity. But one should keep in mind that only $\alpha_{i}$ and $\beta_{i}$ are the independent phase parameters chosen in this work.
Of course, it is always possible to simplify the above linear phase combinations to the forms of $\sin\alpha_{i}$ and $\sin\beta_{i}$ (for $i=1,2,3$) multiplied by the corresponding coefficients, such that the Jarlskog invariant ${\cal J}_{\nu}$ can be compactly expressed as
${\cal J}_{\nu}=\sum^{3}_{i=1}\left(C_{\alpha i}\sin\alpha_{i}+C_{\beta i}\sin\beta_{i}\right)\;,$ (17)
where the coefficients $C_{\alpha i}$ and $C_{\beta i}$ (for $i=1,2,3$) are some formidably complicated algebraic functions of the original seesaw flavor parameters, including $\cos\alpha_{i}$ and $\cos\beta_{i}$ (for $i=1,2,3$). In this way one may actually extract the explicit expressions of $C_{\alpha i}$ and $C_{\beta i}$ from Eqs. (11)—(15), but the relevant results are too lengthy to be presented here.
It is worth pointing out that the 240 phase combinations listed above will be greatly reduced to [33]
$\sin 2\alpha_{1}\;,\;\;\sin 2\alpha_{2}\;,\;\;\sin 2\alpha_{3}\;;\;\;\sin\left(\alpha_{1}\pm\alpha_{2}\right)\;,\;\;\sin\left(\alpha_{2}\pm\alpha_{3}\right)\;,\;\;\sin\left(\alpha_{3}\pm\alpha_{1}\right)\;,$ (18)
if the third heavy Majorana neutrino species $N_{6}$ is switched off, in which case the canonical seesaw mechanism is simplified to the so-called minimal seesaw scenario [34, 35, 36, 37]. Such an amazing simplification of the expression of ${\cal J}_{\nu}$ implies that a lot of information about CP violation in the canonical case is lost in the minimal seesaw framework, although the latter is a good benchmark model.
3 Comments on $\varepsilon_{j\alpha}$
Now let us examine how CP violation in the lepton-number-violating decays of three heavy Majorana neutrinos depends on the original seesaw phase parameters. Far above the electroweak symmetry breaking scale, one may calculate the CP-violating asymmetries between $N_{j}$ decays into the leptonic doublet $\ell_{\alpha}$ plus the Higgs doublet $H$ and their CP-conjugated processes (for $j=4,5,6$ and $\alpha=e,\mu,\tau$) at the one-loop level in the canonical seesaw mechanism [38, 39, 40, 41, 42, 43]. With the help of our Euler-like block parametrization of the seesaw flavor structure, we explicitly obtain
$\varepsilon_{j\alpha}\equiv$ $\hskip 3.69899pt\frac{\Gamma({N}_{j}\to\ell_{\alpha}+H)-\Gamma({N}_{j}\to\overline{\ell_{\alpha}}+\overline{H})}{\displaystyle\sum_{\alpha}\left[\Gamma({N}_{j}\to\ell_{\alpha}+H)+\Gamma({N}_{j}\to\overline{\ell_{\alpha}}+\overline{H})\right]}$
$\simeq$ $\hskip 3.69899pt\frac{1}{\displaystyle 8\pi\langle H\rangle^{2}\sum_{\beta}\left|R_{\beta j}\right|^{2}}\sum^{6}_{k=4}\Bigg{\{}M^{2}_{k}\hskip 2.84544pt{\rm Im}\Bigg{[}\left(R^{*}_{\alpha j}R_{\alpha k}\right)\sum_{\beta}\Big{[}\left(R^{*}_{\beta j}R_{\beta k}\right)\xi(x_{kj})+\left(R_{\beta j}R^{*}_{\beta k}\right)\zeta(x_{kj})\Big{]}\Bigg{]}\Bigg{\}}\;,$ (19)
where the Latin and Greek subscripts run respectively over $(4,5,6)$ and $(e,\mu,\tau)$, $\langle H\rangle\simeq 174~{}{\rm GeV}$ is the vacuum expectation value, $\xi(x_{kj})=\sqrt{x_{kj}}\left\{1+1/\left(1-x_{kj}\right)+\left(1+x_{kj}\right)\ln\left[x_{kj}/\left(1+x_{kj}\right)\right]\right\}$ and $\zeta(x_{kj})=1/\left(1-x_{kj}\right)$ with $x_{kj}\equiv{M}^{2}_{k}/{M}^{2}_{j}$ are the loop functions. Substituting Eq. (6) into Eq. (19), we arrive at the explicit results for the $N_{4}$ decays:
$\varepsilon_{4e}=$ $\hskip 3.69899pt\frac{1}{8\pi\langle H\rangle^{2}I_{44}}\Bigg{\{}M^{2}_{5}s_{14}s_{15}\Bigg{[}\xi(x_{54})\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{1}+\alpha_{i}\right)+\zeta(x_{54})\Big{[}s_{24}s_{25}\sin\left(\alpha_{1}-\alpha_{2}\right)$
$+s_{34}s_{35}\sin\left(\alpha_{1}-\alpha_{3}\right)\Big{]}\Bigg{]}-M^{2}_{6}s_{14}s_{16}\Bigg{[}\xi(x_{64})\sum^{3}_{i=1}s_{i4}s_{i6}\sin\left(\gamma_{1}+\gamma_{i}\right)$
$+\zeta(x_{64})\Big{[}s_{24}s_{26}\sin\left(\gamma_{1}-\gamma_{2}\right)+s_{34}s_{36}\sin\left(\gamma_{1}-\gamma_{3}\right)\Big{]}\Bigg{]}\Bigg{\}}\;,$
$\varepsilon_{4\mu}=$ $\hskip 3.69899pt\frac{1}{8\pi\langle H\rangle^{2}I_{44}}\Bigg{\{}M^{2}_{5}s_{24}s_{25}\Bigg{[}\xi(x_{54})\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{2}+\alpha_{i}\right)+\zeta(x_{54})\Big{[}s_{14}s_{15}\sin\left(\alpha_{2}-\alpha_{1}\right)$
$+s_{34}s_{35}\sin\left(\alpha_{2}-\alpha_{3}\right)\Big{]}\Bigg{]}-M^{2}_{6}s_{24}s_{26}\Bigg{[}\xi(x_{64})\sum^{3}_{i=1}s_{i4}s_{i6}\sin\left(\gamma_{2}+\gamma_{i}\right)$
$+\zeta(x_{64})\Big{[}s_{14}s_{16}\sin\left(\gamma_{2}-\gamma_{1}\right)+s_{34}s_{36}\sin\left(\gamma_{2}-\gamma_{3}\right)\Big{]}\Bigg{]}\Bigg{\}}\;,$
$\varepsilon_{4\tau}=$ $\hskip 3.69899pt\frac{1}{8\pi\langle H\rangle^{2}I_{44}}\Bigg{\{}M^{2}_{5}s_{34}s_{35}\Bigg{[}\xi(x_{54})\sum^{3}_{i=1}s_{i4}s_{i5}\sin\left(\alpha_{3}+\alpha_{i}\right)+\zeta(x_{54})\Big{[}s_{14}s_{15}\sin\left(\alpha_{3}-\alpha_{1}\right)$
$+s_{24}s_{25}\sin\left(\alpha_{3}-\alpha_{2}\right)\Big{]}\Bigg{]}-M^{2}_{6}s_{34}s_{36}\Bigg{[}\xi(x_{64})\sum^{3}_{i=1}s_{i4}s_{i6}\sin\left(\gamma_{3}+\gamma_{i}\right)$
$+\zeta(x_{64})\Big{[}s_{14}s_{16}\sin\left(\gamma_{3}-\gamma_{1}\right)+s_{24}s_{26}\sin\left(\gamma_{3}-\gamma_{2}\right)\Big{]}\Bigg{]}\Bigg{\}}\;.$ (20)
A sum of these three CP-violating asymmetries, denoted as $\varepsilon_{4}\equiv\varepsilon_{4e}+\varepsilon_{4\mu}+\varepsilon_{4\tau}$, is usually referred to as the flavor-independent asymmetry. The expression of $\varepsilon_{4}$ turns out to be independent of the loop functions $\zeta(x_{54})$ and $\zeta(x_{64})$:
$\varepsilon_{4}=$ $\hskip 3.69899pt\frac{1}{8\pi\langle H\rangle^{2}I_{44}}\Bigg{[}\sum^{3}_{i=1}\sum^{3}_{i^{\prime}=1}s_{i4}s_{i^{\prime}4}\Big{[}M^{2}_{5}\xi(x_{54})s_{i5}s_{i^{\prime}5}\sin\left(\alpha_{i}+\alpha_{i^{\prime}}\right)-M^{2}_{6}\xi(x_{64})s_{i6}s_{i^{\prime}6}\sin\left(\gamma_{i}+\gamma_{i^{\prime}}\right)\Big{]}\Bigg{]}\;.$ (21)
A salient feature of the results in Eqs. (20) and (21) is that the phase parameters $\beta_{i}$ (for $i=1,2,3$) do not contribute to CP violation in the $N_{4}$ decay modes.
It is straightforward to work out the CP-violating asymmetries $\varepsilon_{5\alpha}$ and $\varepsilon_{6\alpha}$ (for $\alpha=e,\mu,\tau$) in the same way. We find that the analytical expressions of $\varepsilon_{5\alpha}$ and $\varepsilon_{5}$ can be easily written out from those of $\varepsilon_{4\alpha}$ and $\varepsilon_{4}$ with the subscript replacements $\left(4,5,6\right)\to\left(5,4,6\right)$ together with $\left(\alpha_{i},\gamma_{i}\right)\to-\left(\alpha_{i},\beta_{i}\right)$; and similarly the expressions of $\varepsilon_{6\alpha}$ and $\varepsilon_{6}$ can be directly read off from those of $\varepsilon_{4\alpha}$ and $\varepsilon_{4}$ with the help of the subscript replacements $\left(4,5,6\right)\to\left(6,5,4\right)$ together with $\left(\alpha_{i},\gamma_{i}\right)\to-\left(\beta_{i},\gamma_{i}\right)$.
As a result, CP violation in the decays of three heavy Majorana neutrinos only involve the following 27 linear combinations of the 6 original seesaw phase parameters:
$\begin{array}[]{lllllllll}\sin 2\alpha_{1},&\sin 2\alpha_{2},&\sin 2\alpha_{3};&\sin 2\beta_{1},&\sin 2\beta_{2},&\sin 2\beta_{3};&\sin 2\gamma_{1},&\sin 2\gamma_{2},&\sin 2\gamma_{3};\end{array}$
$\begin{array}[]{llllll}\sin\left(\alpha_{1}\pm\alpha_{2}\right),&\sin\left(\alpha_{2}\pm\alpha_{3}\right),&\sin\left(\alpha_{3}\pm\alpha_{1}\right);&\sin\left(\beta_{1}\pm\beta_{2}\right),&\sin\left(\beta_{2}\pm\beta_{3}\right),&\sin\left(\beta_{3}\pm\beta_{1}\right);\\ \sin\left(\gamma_{1}\pm\gamma_{2}\right),&\sin\left(\gamma_{2}\pm\gamma_{3}\right),&\sin\left(\gamma_{3}\pm\gamma_{1}\right).&&&\end{array}$
One can see that each of the three $\varepsilon_{j}$ asymmetries (for $j=4,5,6$) depends only on 12 simple phase combinations, and each of the nine $\varepsilon_{j\alpha}$ asymmetries (for $j=4,5,6$ and $\alpha=e,\mu,\tau$) is dependent only upon 10 simple phase combinations. It is certainly allowed to compactly express $\varepsilon_{j\alpha}$ and $\varepsilon_{j}$ as
$\varepsilon_{j\alpha}=$ $\hskip 2.84544pt\sum^{3}_{i=1}\left(C^{\prime}_{\alpha i}\sin\alpha_{i}+C^{\prime}_{\beta i}\sin\beta_{i}\right)\;,$
$\varepsilon_{j~{}}=$ $\hskip 2.84544pt\sum^{3}_{i=1}\left(C^{\prime\prime}_{\alpha i}\sin\alpha_{i}+C^{\prime\prime}_{\beta i}\sin\beta_{i}\right)\;,$ (22)
where the coefficients $C^{\prime}_{\alpha i}$, $C^{\prime}_{\beta i}$, $C^{\prime\prime}_{\alpha i}$ and $C^{\prime\prime}_{\beta i}$ (for $i=1,2,3$) can be easily extracted from the expressions of $\varepsilon_{j\alpha}$ and $\varepsilon_{j}$ (for $j=4,5,6$ and $\alpha=e,\mu,\tau$) obtained above. A comparison between Eqs. (17) and (22) indicates that a direct relationship between ${\cal J}_{\nu}$ and $\varepsilon_{j\alpha}$ (or $\varepsilon_{j}$) can in principle be established.
4 Conclusion
It is true that “for neutrino masses, the considerations have always been qualitative, and despite some interesting attempts, there has never been a convincing quantitative model of the neutrino masses”, as remarked by Edward Witten in 2001 [44]. One of the main reasons for this unfortunate situation is that none of the theoretically well-motivated mechanisms, including the elegant canonical seesaw mechanism under discussion, is capable of determining its flavor structure from the very beginning. One possible way out is certainly to invoke possible lepton flavor symmetries [10, 45, 46, 47, 48, 49, 50], but so far none of them stands out as a convincing basis of model building.
In this case we find that a full Euler-like block parametrization of the flavor structure in the seesaw mechanism [11, 12] can help a lot because it makes some important issues analytically calculable. The present work gives a good example of this kind — a general and explicit connection between CP violation in the flavor oscillations of three light neutrino species and that in the lepton-number-violating decays of three heavy Majorana neutrinos is for the first time established via the seesaw bridge. Given that the canonical seesaw mechanism is the most natural, economical and convincing theory for neutrino mass generation and baryogenesis via leptogenesis, we expect that our model-independent result will pave the way for probing the cosmic matter-antimatter asymmetry from CP violation in neutrino oscillations at low energies.
We conclude that there is surely a direct connection between the high- and low-scale effects of CP violation in the seesaw framework, although this seesaw-bridged connection is rather complicated and thus sets a big challenge to its experimental testability.
Acknowledgements
[…]